OpenStudy (anonymous):

³√-216x³ I need help with this problem

OpenStudy (anonymous):

first look at -216*x^3 as (-1)(x^3)*(6^3). Taking the third root you get x*6*thirdroot(-1) -1=i^2 so thirdroot(-1) = thirdroot (i^2) = i^(2/3) so full answer is 6x*i^(2/3)

OpenStudy (anonymous):

the answer will be done like this \[\sqrt[3]{-1}\sqrt[3]{216}\sqrt[3]{x ^{3}}\] which will be -6x i hope u understood